### v2

# Pareto-optimality of pulses for robust population transfer in a ladder-type qutrit

Abstract
Frequency-modulation schemes have recently emerged as alternatives to standard Rabi pulses for realizing robust quantum operations. In this work, we investigate short-duration population transfer between the ground and first excited states of a ladder-type qutrit, with the goal of minimizing leakage into the second excited state. Our multiobjective approach seeks to reduce the maximum transient second-state population and maximize detuning robustness. Inspired by two-state models—such as the Allen-Eberly and Hioe-Carroll models—we extend these concepts to our system, exploring a range of pulse families, including those with super-Gaussian envelopes and polynomial detuning functions. We identify Pareto fronts for pulse models constructed from one of two envelope functions paired with one of four detuning functions. We then analyze how each Pareto-optimal pulse parameter influences the two Pareto objectives and amplitude robustness.
pacs:
Valid PACS appear here
I Introduction
Robust population transfer remains a key area of interest in quantum control. Various techniques employing adiabatic passage have been the primary means to this end, engineered to deliver high fidelity and efficiency with minimal dissipation. The principle behind these methods lies in the concept of adiabaticity, which ensures that a system remains in its instantaneous eigenstate throughout its evolution. For example, in a two-level system with Hamiltonian $\hat{H}=\frac{\hbar}{2}(\Delta\hat{\sigma}_{x}+\Omega\hat{\sigma}_{z})$, the energy eigenvalues are $E_{\pm}=\pm\frac{\hbar}{2}\sqrt{\Delta^{2}+\Omega^{2}}$, corresponding to eigenstates $|-\rangle=\frac{1}{\sqrt{2}}(\cos(\theta/2)|0\rangle-\sin(\theta/2)|1\rangle)$ and $|+\rangle=\frac{1}{\sqrt{2}}(\sin(\theta/2)|0\rangle+\cos(\theta/2)|1\rangle)$, where $\theta$ is the so-called mixing angle and relates to the amplitude and detuning via $\tan(\theta)=\Omega/\Delta$. The system evolves adiabatically if $\dot{\theta}/\Delta E\ll 1$, which is satisfied when $\frac{\norm{\dot{\Omega}\Delta-\Omega\dot{\Delta}}}{(\Omega^{2}+\Delta^{2})^{3/2}}\ll 1$ [1]. When this condition is met, population transfer occurs through gradual evolution along eigenstates, avoiding unwanted transitions.
Rapid adiabatic passage (RAP) is the staple technique for robust population inversion. RAP was first realized in NMR, where the spins on the magnetic moments are inverted by sweeping the drive field frequency across the resonance [2]. It is rapid in the sense that the frequency sweep is faster than the relaxation time of the spins, or in the case of circuit QED, the relaxation time of the qubit. RAP is very versatile in that it can be implemented on any two-level quantum system with an external drive field. As such, it has found many applications in atomic, molecular, and optical physics [3, 4, 5, 6], as well as in chemical reactions [7], interferometry [8, 9], and circuit QED [10, 11]. In circuit QED, the cited works utilize Stark-chirped rapid adiabatic passage (SCRAP), a variety of RAP in which a slowly varying Stark shift sweeps the energy of the target state through resonance to induce complete population transfer [12].
In the context of RAP, pulse shaping is essential for ensuring efficient and robust population transfer. Gaussian pulses are often employed [13, 14], in part because they have been shown to reduce leakage beyond the computational space [15]. Super-Gaussian pulses have also been utilized to enhance control over transition dynamics. For instance, our previous work on high-fidelity qubit control using phase-modulated pulses demonstrated that super-Gaussian envelopes can suppress amplitude more effectively compared to their Gaussian counterparts, providing improved robustness against parameter variations [16]. Since the phase modulation we employ is equivalent to a frequency modulation through the choice of a frame rotation, we hereafter refer the scheme as one that utilizes frequency modulation.
A central goal in quantum control of weakly anharmonic systems, such as the transmon, is to suppress leakage outside the computational subspace—that is, to confine population dynamics to the lowest two energy levels. When driving the ground-to-first-excited-state transition $\mathrm{g-e}$ with a resonant microwave pulse, the close proximity of $\mathrm{e-f}$ transition frequency means that higher levels cannot be ignored. The spectral width of the pulse—particularly if it is short and strong—can unintentionally drive the $\mathrm{e-f}$ transition, leading to population leakage out of the computational subspace. To mitigate this, techniques such as derivative removal by adiabatic gate (DRAG) pulses are used, which shape the drive to cancel unwanted transitions [17]. Alternatively, using longer pulses reduces spectral overlap but can increase sensitivity to decoherence [18]. Ultimately, achieving high-fidelity control in such systems requires carefully balancing gate speed, spectral selectivity, and pulse shaping to minimize leakage while maintaining fast, accurate operations [19, 20].
In this work, we pursue this goal by building on rapid adiabatic passage (RAP) methods and our aforementioned work, where the drive frequency is modulated, in the simplest case, by adding to the qubit frequency an instantaneous linear detuning $\Delta(t)=2\Delta_{\rm{max}}t/T$, parametrized in terms of modulation depth $\Delta_{\rm{max}}$ and pulse duration $T$, with $t\in[-T/2,T/2]$. Furthermore, following our frequency-modulation technique, we employ a single continuously modulated pulse and explore different RAP-like models on a three-level ladder system with states $\{|\rm{g}\rangle,|\rm{e}\rangle,|\rm{f}\rangle\}$, and seek to minimize leakage into the second excited state $|\rm{f}\rangle$ when driving population from the ground state to the first excited state. Specifically, we wish to minimize the maximum transient second excited state population $\max_{t}(p_{\rm{f}})$, with $t\in[-T/2,T/2]$, and simultaneously maximize the detuning robustness at a threshold of the final first excited state population $p_{\rm{e}}(T/2)\geq 0.99$. As this is a multiobjective optimization problem, we seek to find the so-called Pareto front which meets our criteria. A Pareto front is a solution to the multiobjective optimization in which it proves impossible to improve one objective without making the other worse. Pareto fronts have found applications in a plethora of disciplines including economics [21], traffic regulation [22], machine learning [23], and bioinformatics [24]. In physics, Pareto’s methodology has been employed to find the extrema of observables as a form of optimal quantum control [25, 26].
This paper is organized as follows: In Section II we describe the methodology for our frequency-modulated robust population transfer scheme, introduce the pulse models used in this work, and detail the underlying methodology for the Pareto optimization. In Section III, we present the Pareto fronts for the different models, each at two pulse durations, and assign an effective limit to the fronts by utilizing the Landau-Zener formula. We perform parameter analysis in Section IV, and ascertain the most impactful parameters for each of our objective outcomes. Finally, we summarize the key findings of this work in Section V.
II Methodology
We employ a single, continuously modulated drive pulse with Rabi amplitude $\Omega(t)$ and detuning $\Delta(t)$ corresponding to the first transition in order to emulate rapid adiabatic passage between the ground state $|\rm{g}\rangle$ and the first excited state $|\rm{e}\rangle$. For brevity, we hereafter drop the subscript for the Rabi amplitude and detuning function, except when presenting the system Hamiltonian in the next subsection. The shape of the drive pulse plays a central role in achieving efficient and robust state transfer. As we describe in the subsection after next, we focus on two flat-at-the-top envelope functions that are known to deliver strong detuning robustness.
II.1 The frequency-modulation scheme adopted to a ladder qutrit
An energy diagram of the modulation scheme is shown in Fig. 1. The diagonal part of the Hamiltonian in Eq. 1 corresponds to a ladder-type qutrit with energy level separations $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$, in the frame co-rotating with the drive frequency. For a transmon qutrit, the ratio of the Josephson energy $E_{J}$ to the charging energy $E_{C}$ is very large, and in the extreme limit $E_{J}/E_{C}\rightarrow\infty$, the transmon energy levels approach those of a harmonic oscillator for low-energy states, with nearly constant but slightly decreasing spacing due to residual anharmonicity. The qutrit is driven by a microwave field with frequency $\omega_{d}(t)$.
We consider a rotating frame such that for the ground state $|\rm{g}\rangle$ the energy is shifted by $-\Delta(t)$, and the zero of the energy is set at the first excited state $|\rm{e}\rangle$, where $\Delta(t)=\omega_{\rm ge}-\omega_{d}-(d\omega_{d}/dt)t$. The charging energy $E_{C}$ gives rise to the anharmonicity in the energy levels in superconducting qutrits such as the transmon, and relates to the transition frequencies via $E_{\rm C}=\hbar\omega_{\rm{ge}}-\hbar\omega_{\rm{ef}}$. Thus, the detuning of the $|\mathrm{f}\rangle$ level is $\Delta(t)-E_{\rm C}/\hbar$. The ratio of the Rabi frequencies of the two transitions is given by the ratio of the transition electrical dipole moments. The dipole operator is proportional to the number operator $\hat{n}=-i\partial/\partial\varphi$, and thus the dipole matrix elements are $g_{n,n+m}\sim\langle n|\hat{n}|n+m\rangle$, and consequently, $g_{12}/g_{01}=\sqrt{2}$ and $g_{02}=0$.
Combining these considerations, we arrive at the Hamiltonian in the rotating wave approximation, as detailed in Ref. [27]:
$H(t)=\hbar\begin{pmatrix}-\Delta(t)&\Omega(t)/2&0\\ \Omega(t)/2&0&\Omega(t)/\sqrt{2}\\ 0&\Omega(t)/\sqrt{2}&\Delta(t)-E_{C}/\hbar\end{pmatrix}.$ (1)
Figure 1: The energy diagram of our frequency modulation scheme shown for two different detuning functions, one linear (blue), and one quintic (green), for time $t\in[-T/2,T/2]$. The transition frequencies are $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$, the modulation depth is denoted by $\Delta_{\rm max}$, and the frequency offset is $\delta$. The two-photon transition frequency $\omega_{\rm{gf}}^{\rm 2ph}$ is exactly halfway between $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$.
In the following subsection, we discuss the Pareto-optimization technique for finding parameters that, when inserted into the matrix elements of Eq. 1, guide the system to adiabatically follow its lowest eigenstate, thereby facilitating robust population transfer from $|\rm{g}\rangle$ to $|\rm{e}\rangle$.
II.2 Pulse models
Several envelope functions have been explored in two-state and three-state quantum models, each with distinct advantages and limitations [28, 29, 30, 31, 32, 33, 34, 14, 35]. Constant-amplitude pulses can drive transitions via Rabi oscillations but require precise timing and are highly sensitive to detuning errors [36]. Landau-Zener transitions, which use a linearly time-dependent detuning with a constant coupling, offer analytical solutions but can suffer from unwanted population leakage [37, 38]. Further, exponential pulses have been applied in time-dependent quantum transition models but often suffer from long tails, reducing efficiency by causing non-negligible residual population transfer outside the intended time window [39, 40].
Smoothly varying pulses, such as Gaussian and hyperbolic secant envelopes, provide greater control over transition probabilities. Gaussian pulses minimize spectral side lobes and have been widely used in adiabatic passage schemes [13, 41, 42, 43, 14, 35]. However, achieving complete population transfer with Gaussian pulses requires precise detuning control. On the other hand, hyperbolic secant pulses are analytically solvable and naturally suited for robust adiabatic and nonadiabatic population transfer, making them widely used in RAP and related techniques [28, 29, 30, 31, 32, 44, 45].
In this work, we focus on two pulse envelopes: generalized super-Gaussian and hyperbolic secant varieties. Both envelope types are flat at the top, meaning the pulse amplitude remains approximately constant near the center of the drive, around $t=0$, where the avoided crossing typically occurs. This flat-top structure grants a greater degree of detuning robustness by maintaining a high Rabi frequency over the critical transition region, thereby extending the adiabatic window and improving population transfer even under moderate detuning errors [16]. The super-Gaussian envelope allows for greater control over the transition window by flattening the peak and steepening the decay, which can simultaneously be advantageous in terms of requiring lowering amplitudes. Meanwhile, the hyperbolic secant envelope serves as a benchmark due to its well-established analytical properties. These envelope functions are defined as follows:
$\Omega(t)=\Omega_{0}\Bigg{[}\frac{e^{\beta|2t/T|^{n}}-e^{\beta}}{1-e^{\beta}}\Bigg{]}\;,$ (2)
for the super-Gaussian envelope, and
[…]
By employing these two pulse shapes, we aim to explore their effects on state transfer efficiency and robustness, particularly in the context of our frequency-modulation scheme adopted to a three-level system. Each of these envelopes contain three parameters which we seek to multiobjectively optimize with respect to the minimization of the maximum second excited state population $\max_{t}(p_{\rm{f}})$ reached across the pulse duration $t\in[-T/2,T/2]$, and (simultaneously) with respect to the detuning robustness defined as the frequency width of the first excited state population at the threshold $p_{\rm{e}}(T/2)\geq 0.99$ with a local maximum of $p_{\rm{e}}(T/2)\geq 0.999$ within this width. The envelope parameters are amplitude $\Omega_{0}$, envelope order $n$, and the parameter $\beta$ (for super-Gaussian pulses) or $\epsilon$ (for hyperbolic secant pulses), which determine the envelope width and steepness. The variable order $n$ generalizes the super-Gaussian envelopes in that if $n=2$ we recover a Gaussian profile, $n=4$ corresponds to the so-called super-Gaussian variety, and as $n\rightarrow\infty$ the pulse approaches a rectangular form. Amplitudes $\Omega_{0}$ are searched for in the range $[0,0.1]$ GHz, consistent with experimental constraints that help mitigate parasitic effects such as leakage into $|\rm{f}\rangle$ and higher states, as well as AC Stark shifts, which are known to degrade qubit coherence and fidelity in similar systems [46]. The envelope order $n$ is scanned over the range $(0,10]$, while $\beta$ and $\epsilon$ are searched within $[-6,0)$ and $(0,6]$, respectively.
Additionally, we consider four detuning models and assign each to the two envelope types. Our first detuning model is linear, and the model name when assigned to the super-Gaussian envelope is SG1, and SECH1 when assigned to the hyperbolic secant envelope. For subsequent models, we hereafter adopt the prefixes SG and SECH for the model names. The modulation depth $\Delta_{\rm max}$ has an upper limit of half the charging energy, which is set at $E_{C}/2=0.15$ GHz in this work, since it is the maximum depth over which we can sweep in a typical circuit QED experiment using a transmon.
The first detuning model has the form
$\Delta(t)=\frac{2k_{1}}{T}t+\delta\;,$ (4)
where the prefactor $k_{1}$ determines the slope of our detuning function and is a parameter we optimize by searching in the range $k_{1}\in[-0.3,0.3]$ GHz. Since Eq. 4 is a linear detuning function, the parameter $k_{1}=\Delta_{\rm{max}}$. Further, we center our modulation about the frequency offset $\delta$, i.e., by sweeping $(2\Delta_{\rm max}/T)t$ from $t=-T/2$ to $t=T/2$.
We also explore a quintic detuning function, i.e.,
$\Delta(t)=\sum\limits_{n=1}^{5}k_{n}(2t/T)^{n}+\delta,$ (5)
where $k_{n>1}$ are each searched for in the range $[-0.1,0.1]$ GHz. For this detuning function, the effect of the parameter $k_{1}$ is the same as in Eq. 4; whereas, the prefactor for the quadratic term, $k_{2}$, makes the detuning function convex (upward-bending) when $k_{2}>0$ and concave (downward-bending) when $k_{2}<0$. Further, the parameter $k_{3}$, the cubic prefactor, introduces asymmetry and has the effect of making the right side ($t>0$) of the detuning function steeper and the left side ($t<0$) shallower when $k_{3}>0$, and vice versa when $k_{3}<0$. Still further, the quartic ($k_{4}$) and quintic ($k_{5}$) prefactors have similar effects to $k_{2}$ and $k_{3}$, respectively, but with stronger effects at larger |$\absolutevalue{t}$ |. Overall, the symmetry properties are governed by the parity of each term: the even-powered terms ($k_{2}$, $k_{4}$) contribute symmetric components about $t=0$, while the odd-powered terms ($k_{1}$, $k_{3}$, $k_{5}$) contribute antisymmetric components. A detuning function composed purely of odd-powered terms is antisymmetric about the origin, but adding even-powered terms breaks this antisymmetry, resulting in an overall asymmetric function. Thus, the inclusion of both even and odd terms enables finer control over both the shape and symmetry of the detuning profile.
We also examine detuning functions that incorporate the hyperbolic tangent. Motivated by the Allen–Eberly [31] and Hioe–Carroll [32] models—both of which employ hyperbolic secant pulse envelopes—we respectively consider the following forms:
$\Delta(t)=k_{1}\tanh(2\gamma_{1}t/T)+\delta,$ (6)
and
[…]
where we subtract by $k_{2}$ in Eq. 7 to ensure that we satisfy the condition $\Delta(0)-\delta=0$, which is naturally satisfied for the other detuning functions we consider. For these detuning functions, the parameters $k_{1}$ and $k_{2}$ are searched for in the aforementioned ranges $k_{1}\in[-0.3,0.3]$ GHz and $k_{2}\in[-0.1,0.1]$ GHz; whereas, the parameters $\gamma_{1}$ and $\gamma_{2}$ are each searched for in the range $(0,8]$. The parameter $k_{1}$ acts as a scaling factor which symmetrically affects at which value the detuning saturates, i.e., the detuning asymptotically approaches ±$\pm k_{1}$ since $\tanh(x)\rightarrow\pm 1$ when |$\absolutevalue{x}$ | is large. However, $\gamma_{1}$ controls how quickly the detuning transitions from negative to positive values, i.e., the steepness of the slope. Small values of $\gamma_{1}$ yield a near-linear detuning and large values yield a sharper transition near $t=0$, with extremely large values producing a step-like detuning. Since Eq. 7 is just Eq. 6 plus an additional term, the effect of these parameters is the same in both detuning functions. Moreover, the additional term in Eq. 7 introduces a dip in the detuning function around $t=0$ and increasing $k_{2}$ makes the dip more pronounced. The parameter $\gamma_{2}$ affects the localization of the dip, with a larger $\gamma_{2}$ yielding a dip more concentrated around $t=0$ and a smaller $\gamma_{2}$ making it spread out over a longer time.
As we have previously introduced the shorthand notations SG and SECH for our two envelope types, we further systematically distinguish our models by referring to Eqs. 4, 5, 6, and 7 as detunings of type 1, type 2, type 3, and type 4, respectively. We then create our model names by agglutinating the SG or SECH suffix to a particular detuning model as shown in Table 1. Thus, in this work, we consider a total of eight models: four with super-Gaussian envelopes and four with hyperbolic secant envelopes. These models each satisfy the conditions $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, $\Delta(0)=\delta$, and are evaluated at pulse durations of $T=50$ ns and $T=200$ ns. These two pulse durations span a range of experimentally relevant gate times in superconducting qubit hardware. Shorter durations are motivated by the need for fast gate operations in near-term devices, whereas longer durations allow greater detuning robustness and reduced leakage into higher excited states. As an example, in transmon-based systems, we have previously shown that sufficiently long pulses driven close to resonance with modest modulation depths can exhibit detuning robustness with a negligible final second excited state population, thereby isolating the qubit subspace despite the transmon’s weak anharmonicity [16].
Two-state-inspired models
Model Envelope $[\Omega(t)/\Omega_{0}]$ Detuning $[\Delta(t)-\delta]$
SG1 $\left(\frac{e^{\beta|2t/T|^{n}}-e^{\beta}}{1-e^{\beta}}\right)$ $2k_{1}t/T$
SG2 $\left(\frac{e^{\beta|2t/T|^{n}}-e^{\beta}}{1-e^{\beta}}\right)$ $\sum\limits_{n=1}^{5}k_{n}(2t/T)^{n}$
SG3 $\left(\frac{e^{\beta|2t/T|^{n}}-e^{\beta}}{1-e^{\beta}}\right)$ $k_{1}\tanh(2\gamma_{1}t/T)$
SG4 $\left(\frac{e^{\beta|2t/T|^{n}}-e^{\beta}}{1-e^{\beta}}\right)$ $k_{1}\tanh(2\gamma_{1}t/T)$
[…]
Table 1: The two-state-inspired models used in this work, each consisting of an amplitude (pulse envelope) function $\Omega(t)$ and detuning function $\Delta(t)$. Each envelope function has a total of three parameters, i.e., $[\Omega_{0},n,\beta\;(\epsilon)]$ for SG (SECH) envelopes. Here, the detuning prefactors are denoted as $k_{1}$, $\cdots$, $k_{5}$, with detuning models of type 1 and 3 only using $k_{1}$, type 4 using $k_{1}$ and $k_{2}$, and type 2 using all 5. The parameters $\gamma_{1}$, and $\gamma_{2}$ only appear in the detuning functions of types 3 and 4. Each model satisfies the conditions $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, and $\Delta(0)=\delta$.
II.3 Pareto-optimization
The conundrum of determining the optimal pulse parameters for our models arises from the fact that a pulse that maximizes detuning robustness $\Delta_{\rm{rob}}$ at a given threshold of $p_{\rm{e}}(T/2)$ will generally not minimize the maximum transient second-excited state population, $\max_{t}(p_{\rm{f}})$, and vice versa. This creates a multiobjective optimization (MOO) problem, where we must balance these competing objectives. To address this trade-off, we apply evolutionary computation methods to identify the best solutions. Techniques such as genetic algorithms and estimation of distribution algorithms have been successfully used in complex optimization problems across various fields, including evacuation planning [47] and robotic control optimization [48]. Genetic algorithms operate by mimicking the principles of natural selection to iteratively evolve a population of candidate solutions toward optimality. The process begins with the generation of an initial population that satisfies predefined bounds and constraints. Each individual in the population is evaluated using an objective function to assign a fitness score. The algorithm then progresses through iterative stages, where parents for the next generation are selected based on their fitness [49].
In solving this MOO, we rely on the Pareto front, which represents the optimal trade-off between objectives in the objective space. Pareto fronts have been used in multiobservable quantum control, where it was shown that the smoothness and convexity of the Pareto front depend on the commutativity of the observables [26]. To find the Pareto front numerically, we use MATLAB’s gamultiobj() function, which implements a controlled, elitist genetic algorithm based on the non-dominated sorting genetic algorithm (NSGA-II) [50].
Building on these general principles, NSGA-II enhances the process with additional layers of sophistication, such as non-dominated sorting and crowding distance measures. After scoring the offspring based on their objective function values and feasibility, the current population and offspring are merged into a single set. This set is sorted into non-dominated fronts and individuals are ranked based on both their Pareto rank and their crowding distance—a measure of solution diversity within a front. Finally, the population is trimmed to its original size, retaining individuals from higher-ranking fronts and preserving diversity. This iterative process continues until convergence criteria are met, yielding an accurate approximation of the Pareto front and enabling efficient exploration of multiobjective optimization landscapes.
III Results
In this section, we present the Pareto fronts for each of our models which optimally minimizes the maximum transient second-state population $\max_{t}(p_{\rm{f}})$ in the range $t\in[-T/2,T/2]$ given that we must simultaneously maximize the detuning robustness at the threshold $p_{\rm{e}}(T/2)\geq 0.99$. These fronts are evaluated for pulse durations of $T=50$ ns and $T=200$ ns. Since this threshold may be reached at multiple points across the detuning (or amplitude) axis, a consistent definition of robustness is essential. We therefore define detuning robustness (or amplitude robustness) as the width of the $p_{\rm{e}}$ curve between the first two points where $p_{\rm{e}}(T/2)=0.99$, provided that $p_{\rm{e}}(T/2)\geq 0.999$ is reached at least once within this interval. We include amplitude robustness in parenthesis, because, even though it was not one of our objectives in our MOO problem, we compute it using the optimal pulse parameters found by the genetic algorithm. However, we do not discuss it until the next section.
Moreover, the Pareto fronts for each of our models at the two pulse durations $T$ are presented in Fig. 2. Here, we observe that the models form two distinct clusters: one for $T=50$ ns pulses (solid lines) and one for $T=200$ ns pulses, with the latter affording us more detuning robustness. For detuning robustness values up to approximately $0.1$ GHz for 50 ns pulses and $0.15$ GHz for 200 ns pulses, all models yield similar $\max_{t}(p_{\rm{f}})$, with the clusters corresponding to respective pulse durations. This suggests that, within these moderate robustness regimes, performance is largely insensitive to the specific choice of pulse model.
We further note that linear detuning functions perform less optimally compared to the other detuning types at higher detuning robustness levels, as evidenced the divergence towards higher $\max_{t}(p_{\rm{f}})$ values at approximately 0.13 GHz for $50$ ns pulses and at approximately 0.22 GHz for $200$ ns pulses. Otherwise, it is generally less obvious how the other models fare relative to each other. Thus, to better contextualize these results, we assign approximate theoretical bounds to these fronts to ascertain how close we are to the limits imposed by adiabaticity.
Figure 2: The maximum transient second excited state population $\max_{t}(p_{\rm{f}})$ vs detuning robustness for various models with pulse durations of 50 ns (solid curves) and 200 ns (dashed curves). The shaded regions (cyan and magenta for 50 ns and 200 ns pulses, respectively) represent the range of data (convex hull) constituting the fronts found from using the Landau-Zener (LZ) formula to approximate the theoretical limit of the detuning robustness for each model. These LZ-based bounds are computed at the threshold condition $p_{\rm{e}}^{\rm{thresh}}=0.99$, but without requiring that $p_{\rm{e}}(T/2)\geq 0.999$ anywhere within the corresponding detuning interval (i.e., between the roots $\delta_{-}$ and $\delta_{+}$, where $p_{\rm{e}}(\delta_{\pm})=0.99$). The orange dotted boundary includes this additional constraint. For the models with nonlinear detuning, $\Delta(t)$ were expanded about $t=0$ and the chirp prefactor (Landau-Zener velocity) of the linear term was used in the LZ formula.
To this end, we use the Landau-Zener (LZ) formula in calculating the detuning robustness at the threshold $p_{\rm{e}}(T/2)\geq 0.99$. The LZ formula provides the probability of a diabatic transition, i.e.,
$p_{\rm LZ}=e^{-2\pi\Gamma},$ (8)
where $\Gamma=\Omega(t_{0})^{2}/4\hbar|v|$, which quantifies how adiabatic the transition is. Here $t_{0}$ is defined as the time at $\Delta(t)=0$, i.e., $t_{0}=-\delta/v$. In the limit $\Gamma\gg 1$, the transition is adiabatic and if $\Gamma\ll 1$, the transition is nonadiabatic.
The chirp prefactor $v$ of the linear term of the detuning function is referred to as the Landau-Zener velocity and represents the slope of $\Delta(t)$ at the avoided crossing, i.e.,
$v=\left.\frac{d\Delta(t)}{dt}\right|_{t=0}.$ (9)
Importantly, $v$ dominates regardless of the detuning function. Specifically, for detunings of type 1 and 2, $v=2k_{1}/T$, whereas for detunings of type 3 and 4, $v=2k_{1}\gamma_{1}/T$ In terms of $v$ and considering the threshold value of the first excited state population we use for calculating the detuning robustness, i.e., $p_{\rm{e}}^{\rm{thresh}}=0.99$, we impose the condition
$1-p_{\rm LZ}\geq p_{\rm{e}}^{\rm{thresh}}\;.$ (10)
Here, we subtract from unity as Eq. 8 describes the probability of remaining in $|\rm{g}\rangle$ at the level crossing, and we further assume that $|\rm{f}\rangle$ is far detuned from the other states (e.g., for $T=200$ ns pulses, $|\rm{f}\rangle$ is only populated if $\delta\approx E_{C}/2$ [16]). Thus, after substituting $t_{0}$ in Eq. 10 such that it explicitly depends on $\delta$, we solve for $\delta$ to find the two roots of the threshold condition $p_{\rm{e}}^{\rm{thresh}}=0.99$, denoted $\delta_{-}$ and $\delta_{+}$. Due to symmetry, $\delta_{-}=-\delta_{+}$, and we approximate the detuning robustness as $\Delta_{\rm{rob}}^{\rm{LZ}}=2\delta_{+}$. Combining $\Delta_{\rm{rob}}^{\rm{LZ}}$ with the previously found values of $\max_{t}(p_{\rm{f}})$, we obtain the bounds in Fig. 2: the cyan-colored region for the 50 ns pulses and the magenta-colored region for the 200 ns pulses. These bounds are less stringent than our definition of detuning robustness, as they do not include the constraint that $p_{\rm{e}}(T/2)\geq 0.999$ anywhere between $\delta_{-}$ and $\delta_{+}$. The corresponding limit that does incorporate this additional constraint is shown as the orange dotted boundary, plotted only for the 50 ns pulses, since for the $200$ ns pulses, the limits with and without this constraint nearly coincide. Notably, the region where the orange dotted boundary does not overlap with the cyan-colored boundary indicates that, under the Landau–Zener approximation, the $50$ ns pulses satisfy our stricter definition of detuning robustness only at higher values of $\max_{t}(p_{\rm{f}})$. We note that all models except for those with linear detunings (type-1) and the SECH2 evaluated for $T=50$ ns, which prove less optimal for part of their fronts, lie firmly within these bounds.
Moreover, Fig. 2 reveals that all pulse models are capable of achieving moderate detuning robustness—for instance, $\Delta_{\rm{rob}}=0.1$ GHz—even when the maximum transient population of the second excited state remains low (i.e., $\max_{t}(p_{\rm{f}})$ $\leq 0.01$). Models featuring nonlinear detuning profiles (types 2–4) become essential only when targeting higher levels of detuning robustness. Shorter pulses ($T=50$ ns) can attain modest robustness but typically incur higher maximum transient second-excited state populations. By contrast, comparable—or in some cases greater—detuning robustness can be achieved with reduced $\max_{t}(p_{\rm{f}})$, albeit at the cost of longer pulse durations ($T=200$ ns).
While the Pareto fronts serve as our primary figure of merit, they tell us nothing explicitly about the optimality of the parameters necessary to create them. In fact, each envelope function has a total of three parameters, i.e., $[\Omega_{0},n,\beta\;(\epsilon)]$ for SG (SECH) pulses, and depending on the detuning function, there are anywhere from one (for type-1) to five (type-2) detuning parameters. In the next section, we ascertain the overall importance of each of these parameters with respect to $\max_{t}(p_{\rm{f}})$, detuning robustness $\Delta_{\rm{rob}}$, and amplitude robustness $\Omega_{\rm{rob}}$.
IV Parameter Analysis
Figure 3: The parameter sensitivity of (a) $\max_{t}(p_{\rm{f}})$, (b) detuning robustness $\Delta_{\rm{rob}}$, and (c) amplitude robustness $\Omega_{\rm{rob}}$, where the vertical axis is the standard deviation $\sigma$ and the horizontal axis is the mean absolute effect $\mu^{*}$ found via the Morris method with a $5\times 10^{-3}\%$ perturbative decrease on each model parameter. The legend is shown in (d).
To assess the sensitivity of the model parameters on each objective outcome, we employ a modified version of the Morris sensitivity test [51]. This method involves calculating two key metrics: the mean absolute effect $\mu^{*}$ and the standard deviation $\sigma$.
The elementary effect for a given parameter is calculated by perturbing the parameter by a fixed percentage (in our case, a $5\times 10^{-3}\%$ decrease) while keeping all other parameters fixed. The elementary effect is then defined as the difference between the perturbed and unperturbed model output, divided by the perturbation step size ($-5\times 10^{-5}$). Unlike the conventional Morris method, we use the relative elementary effect, which is obtained by dividing each elementary effect by the unperturbed output. The mean absolute effect is the average of the absolute values of these relative elementary effects, providing a measure of the overall importance of each parameter in influencing the model output. Explicitly, the mean absolute effect is
$\mu^{*}=\frac{1}{N}\sum\limits_{i=1}^{N}|E_{i}|,$ (11)
where $E_{i}$ is the relative elementary effect for the $i$-th sample of a parameter and $N$ is the total number of samples (175 for each of our parameters).
The standard deviation $\sigma$ measures the variability in these relative elementary effects across the parameter space. This quantifies the degree to which interactions between parameters influence the model output. Specifically, we calculate the sample standard deviation of the relative elementary effects for each parameter, i.e.,
$\sigma=\sqrt{\frac{1}{N-1}\sum\limits_{i=1}^{N}(E_{i}-\mu)^{2}},$ (12)
where $\mu$ is the mean effect and the $N-1$ in the denominator accounts for the potential variability beyond the analyzed set. Unlike $\mu^{*}$, which is the mean of the absolute values of the relative elementary effects (Eq. 11), $\mu$ considers the raw values and retains their signs. The mean effect is a useful metric in its own right, as it indicates whether a parameter’s influence on the output is typically positive or negative, revealing the directionality of its impact. Moreover, the ratios of $\mu/\mu^{*}$, which assume values between $-1$ and $1$, inform us of the type of correlation of the parameters relative to the outputs. We have calculated these ratios for $\max_{t}(p_{\rm{f}})$, detuning robustness, and amplitude robustness for the case of a $5\times 10^{-3}\%$ perturbative decrease on each parameter. For instance, we see that for $\max_{t}(p_{\rm{f}})$, the amplitude $\Omega_{0}$ and envelope order $n$ parameters, each have a completely positive correlation with $\max_{t}(p_{\rm{f}})$ for all models, i.e., $\mu=\mu^{*}$, so the increase in each of these parameters, keeping all others fixed, yields a higher $\max_{t}(p_{\rm{f}})$. The same can be said for the detuning parameter $k_{1}$, except for models of type-4 detuning, i.e., SG4 and SECH4 models. For detuning and amplitude robustness, the ratios $\mu/\mu^{*}$ largely lie between $-1$ and $1$ for the models, which suggests a more complex correlation. On the other hand, the envelope parameters $\beta$ and $\epsilon$ are completely negatively correlated, i.e., $\mu=-\mu^{*}$, for all models when the output is $\max_{t}(p_{\rm{f}})$, and nearly completely negatively correlated when the output is detuning robustness. However, for amplitude robustness, these parameters are generally strongly positively correlated for the models.
The standard deviation quantifies interactions among parameters, as the effect of changing a parameter will vary based on the levels of other parameters if strong interactions are present. This variability causes a higher spread in $E_{i}$, resulting in a larger $\sigma$. Moreover, if increasing one parameter has a different impact depending on whether another parameter is high or low, this interaction will increase the variability of $E_{i}$ for the former parameter, and consequently, $\sigma$ for that parameter will be large.
While $\mu^{*}$ measures the overall importance of a parameter (how much it influences the output on average), $\sigma$ highlights whether that influence is consistent or varies significantly. A parameter with high $\mu^{*}$ but low $\sigma$ suggests a dominant, consistent effect, while high $\mu^{*}$ and high $\sigma$ suggest strong influence with significant variability, likely due to interactions or nonlinearity.
As can be seen in Fig. 3 (a), there is a cluster of $\Omega_{0}$ parameters with high $\mu^{*}$ and low $\sigma$. This cluster solely contains $\Omega_{0}$ for each model and pulse duration, which clearly indicates $\Omega_{0}$ has the most profound effect on the outcome of minimizing $\max_{t}(p_{\rm{f}})$. In Fig. 3 (b) and (c) the standard deviation is generally much higher among the parameters.
V Discussion and conclusions
Inspired by exactly solvable two-state models, we have investigated the Pareto-optimality of population transfer from the ground state to the first excited state in a ladder-type qutrit. Our optimization considered two competing objectives: minimizing the maximum transient population of the second excited state, and maximizing detuning robustness. We define the latter as the width of the first excited state population $p_{\rm{e}}$ curve between the first two points where $p_{\rm{e}}(T/2)=0.99$, provided that $p_{\rm{e}}(T/2)\geq 0.999$ is reached at least once within this interval. Due to the inherent tradeoff between these objectives, the Pareto front represents the set of optimal solutions to this multiobjective problem.
To explore the optimal pulse design, we analyzed eight pulse models, each defined by a combination of an envelope function and a detuning function. The envelope function is either a super-Gaussian or a hyperbolic secant, while the detuning function follows one of four forms: linear (type-1), quintic (type-2), hyperbolic tangent (type-3), or a superposition of hyperbolic secant and hyperbolic tangent (type-4). Each model satisfies the conditions: $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, and $\Delta(0)-\delta=0$. We further examined pulse durations of 50 ns and 200 ns and determined the Pareto-optimal model parameters that achieve the best tradeoff between our objectives.
As shown in Fig. 2, the resulting Pareto fronts cluster into two distinct regions, with the 200 ns pulses yielding more optimal results. To further contextualize these findings, we approximate theoretical bounds using the Landau-Zener formula. By expanding the nonlinear detuning functions around $t=0$ and extracting the linear term (Landau-Zener velocity), we compare our results against this theoretical benchmark. We find that the Pareto fronts—particularly those corresponding to type-3 and type-4 detunings—closely approach the Landau-Zener boundary. Additionally, our analysis reveals that the choice of detuning function has a more significant impact on optimality than choosing between the two flat-at-the-top envelope functions.
Further, all pulse models considered are capable of delivering moderate detuning robustness (e.g., $\Delta_{\rm{rob}}=0.1$ GHz) while keeping the maximum transient population of the second excited state low (i.e., $\max_{t}(p_{\rm{f}})$ $\leq 0.01$). However, achieving higher detuning robustness necessitates the use of nonlinear detuning profiles (types 2–4). Shorter pulses ($T=50$ ns) generally need higher $\max_{t}(p_{\rm{f}})$ to achieve similar robustness, while longer pulses ($T=200$ ns) can reach comparable or even better robustness with lower transient excitation.
Moreover, this trend that all models perform comparably well at moderate detuning robustness but diverge at higher values— is reflected in the structure of the Pareto fronts. For detuning robustness approximately up to 0.1 GHz (50 ns pulses) and 0.15 GHz (200 ns pulses), all models exhibit comparable $\max_{t}(p_{\rm{f}})$, as their fronts cluster tightly into two distinct groups corresponding to the pulse durations. This suggests that within these ranges, performance is largely model-agnostic.
To assess the sensitivity of each model parameter, we employed the Morris method, perturbing each parameter individually and evaluating its effect on $\max_{t}(p_{\rm{f}})$, $\Delta_{\rm{rob}}$, and $\Omega_{\rm{rob}}$. The relative elemental effect is computed as the difference between the perturbed and unperturbed outputs divided by the signed step size and further divided by the unperturbed output. From this, we compute the mean effect ($\mu$), mean absolute effect ($\mu^{*}$), and standard deviation ($\sigma$) of each parameter’s influence. By plotting $\sigma$ versus $\mu^{*}$ (Fig. 3), we identify key trends: For $\max_{t}(p_{\rm{f}})$, the amplitude $\Omega_{0}$ stands out as the most impactful parameter, forming an island of high $\mu^{*}$ and low $\sigma$ across all models. For $\Delta_{\rm{rob}}$ and $\Omega_{\rm{rob}}$, no single parameter dominates, suggesting a more intricate interplay between multiple parameters. However, some exceptions exist, such as $\Omega_{0}$, which has no impact on amplitude robustness.
In conclusion, our study provides a systematic approach for designing optimal pulses that enable robust population transfer in ladder-type qutrits. The insights gained here are broadly applicable to any system where the pulse models discussed can be implemented.
Acknowledgements.
This project has received funding from the EU Flagship on Quantum Technology through HORIZON-CL4-2022-QUANTUM-01-SGA Project No. 101113946 OpenSuperQPlus100.   We also acknowledge support from Research Council of Finland Proof-of-Concept program (Project 359397) and from the Centre of Excellence program (Project 352925).